zarasht:

What is the mass of 5.00 L of propane vapor (C₃H₈) at STP?

Nnesha:

At STP condition: 1 mole of a gas= 22.4 L volume Given: \[ \rm 5.00L~~of~~C_3H_8 \]first we need to get rid of "L" how would we do that? we know 1 mole of a gas = 22.4 L so \[ \rm 5.00L~~C_3H_8 \cdot \frac{ 1~mole~C_3H_8}{22.4~ L~ C_3H_8} \]

Nnesha:

\[ \rm 5.00 \cancel{L~~C_3H_8} \cdot \frac{ 1~mole~C_3H_8}{\cancel{22.4~ L~ C_3H_8}} \] L cancels out, left with \[\rm 1 ~mole~C_3H_8\] do you know how to convert moles to mass ??

Nnesha:

1 mole of C₃H₈ = how many grams of C₃H₈?